Timberly Construction negotiates a lump-sum purchase of several assets from a company that is going out of business. The purchase is completed on January 1, 2017, at a total cash price of $850,000 for a building, land, land improvements, and four vehicles. The estimated market values of the assets are building, $471,750; land, $286,750; land improvements, $46,250; and four vehicles, $120,250. The company’s fiscal year ends on December 31.

Required:

1-a. Prepare a table to allocate the lump-sum purchase price to the separate assets purchased.
1-b. Prepare the journal entry to record the purchase.
2. Compute the depreciation expense for year 2017 on the building using the straight-line method, assuming a 15-year life and a $29,000 salvage value.
3. Compute the depreciation expense for year 2017 on the land improvements assuming a five-year life and double-declining-balance depreciation.
Answer #1

1a) Allocation of lump sum purchase price

Assets Market values Weight Allocation of lump sum price
Building 471750 0.51 850000*.51 = 433500
Land 286750 0.31 850000*.31 = 263500
Land improvement 46250 0.05 850000*.05 = 42500
Vehicles 120250 0.13 850000*.13 = 110500
Total 925000 850000

1b) Journal entry

Date account and explanation debit credit
Building 433500
Land 263500
Land improvement 42500
Four vehicle 110500
Cash 850000
(To record purchase)

2) Depreciation expense for building = (433500-29000/15) = 26967

3) Depreciation expenses for land improvement = 42500*40% = 17000